OpenStudy (swissgirl):

Integrate \(e^{\frac{x}{100}}Sin( \frac{x}{4})\)dx

OpenStudy (swissgirl):

I am getting smth really ugly and idk I was hoping for smth nicer

OpenStudy (kinggeorge):

Well in general, \[\int e^{ax}\sin(bx)\,dx=\frac{e^{ax}}{a^2+b^2}(a\sin(bx)-b\cos(bx))\]So you'll probably end up with something ugly looking.

OpenStudy (swissgirl):

Okkkk I guess sooooo :) Its gonna be a messy word problem :(

OpenStudy (swissgirl):

Thanks :)

OpenStudy (kinggeorge):

You're welcome.